Given the following data, estimate the value of f(5) using
i) Lagrange’s interpolation
ii) Newton’s divided difference interpolation
x 1 4 6 7 10
f(x) -3 9 10 9 8
1
Expert's answer
2020-04-20T11:44:35-0400

"\begin{matrix}
i&0&1&2&3&4\\
 x & 1&4&6&7&10 \\
 f(x)&-3&9&10&9&8
\end{matrix}"

i) Lagrange’s interpolation 

"f(x)=\frac{(x-4)(x-6)(x-7)(x-10)}{(1-4)(1-6)(1-7)(1-10)}\cdot(-3)+\\
+\frac{(x-1)(x-6)(x-7)(x-10)}{(4-1)(4-6)(4-7)(4-10)}\cdot9+\\
+\frac{(x-1)(x-4)(x-7)(x-10)}{(6-1)(6-4)(6-7)(6-10)}\cdot10+\\
+\frac{(x-1)(x-4)(x-6)(x-10)}{(7-1)(7-4)(7-6)(7-10)}\cdot9+\\
+\frac{(x-1)(x-4)(x-6)(x-7)}{(10-1)(10-4)(10-6)(10-7)}\cdot8"

"f(5)=\frac{(5-4)(5-6)(5-7)(5-10)}{(-270)}+\\
+\frac{(5-1)(5-6)(5-7)(5-10)}{(-12)}+\\
+\frac{(5-1)(5-4)(5-7)(5-10)}{(4)}+\\
+\frac{(5-1)(5-4)(5-6)(5-10)}{(-6)}+\\
+\frac{(5-1)(5-4)(5-6)(5-7)}{81}=\\
=\frac{1}{27}+\frac{10}{3}+10-\frac{10}{3}+\frac{8}{81}=10.14"


ii) Newton’s divided difference interpolation  

"\begin{matrix}
 1 & -3 \\
&&\frac{9+3}{4-1}=4\\
 4 & 9&&\frac{0.5-4}{6-1}=-0.7\\
&&\frac{10-9}{6-4}=0.5&&\\
6&10&&\frac{-1-0.5}{1-7}=0.25&&\\
&&\frac{9-10}{7-6}=-1&&\\
7&9&&\frac{-0.3+1}{4-10}=-0.12\\
&&\frac{8-9}{10-7}=-0.3\\
10&8
\end{matrix}"



"\begin{matrix}
 \frac{0.25+0.7}{7-1}=0.16\\
&&\frac{-0.06-0.16}{10-1}=-0.02\\
\frac{-0.12-0.25}{10-4}=-0.06\\
\end{matrix}"

Resulting coefficients are:[-3,4,-0.7,0.16,-0.02]

Create Target Polynomials

"f(x)=-3+4(x-1)-0.7(x-1)(x-4)+\\+0.16(x-1)(x-4)(x-6)-\\
-0.02(x-1)(x-4)(x-6)(x-7)\\
f(5)=-3+4(5-1)-0.7(5-1)(5-4)+\\+0.16(5-1)(5-4)(5-6)-\\
-0.02(5-1)(5-4)(5-6)(5-7)=9.4"